Suppose the monthly demand for soda by a consumer is given by Q = 40 - 10P. At a price of $2 and quantity of 20, what is the own-price elasticity of demand for soda?


1
Expert's answer
2021-03-08T09:06:18-0500

"\bold {Answer}"

Own-price elasticity of demand, "\eta_P = \bold {-1}"

Elasticity is unitary.


"\bold {Solution}"

Own price elasticity of demand "(\eta_p)" is a measure of the sensitivity of quantity demanded to changes in own price of a good.

"\eta_p = \dfrac {dQ}{dP} × \dfrac {P}{Q}"

"\dfrac {dQ}{dP} = \dfrac {d}{dP}(40-10P)"


"= -10"


"When \space P= \$2, \space Q = 20 \space units"


"Thus,"

"\eta_p = -10 × \dfrac {2}{20}"

"= \bold {-1}"


Hence, the own price elasticity of demand is unitary.